OpenStudy (kewlgeek555):

The first five terms of a sequence are shown below. 5, 9, 13, 17, 21 If the nth term of this sequence is represented by f(n), which of the following functions best represents this sequence? f(n) = 5n - 1; n ≥ 1 f(n) = 5n + 1; n ≥ 1 f(n) = 4n - 1; n ≥ 1 f(n) = 4n + 1; n ≥ 1

OpenStudy (kewlgeek555):

@Loser66 I think D.

OpenStudy (kewlgeek555):

I think it is D because it is adding four every time...

OpenStudy (kewlgeek555):

I didn't understand - sorry.

OpenStudy (kewlgeek555):

...

OpenStudy (ranga):

This is an arithmetic progression, with the first term 5 and common difference: 4. The nth term of the sequence: a, a + d, a + 2d, a + 3d, ..... is: a + (n-1)d

OpenStudy (kewlgeek555):

Okay so the first term would be a?

OpenStudy (ranga):

Yes. The first term is a. Here a = 5. The common difference is d. Here d = 4. Put it in the formula for the nth term and simplify.

OpenStudy (kewlgeek555):

So the answer would be A?

OpenStudy (ranga):

No. Put a = 5, d = 4 in a + (n-1)d and simplify.

OpenStudy (kewlgeek555):

5 + (n - 1)4 5 + 4n - 1

OpenStudy (ranga):

No. 5 + (n - 1)4 = 5 + 4n - 4 = 4n + 1

OpenStudy (kewlgeek555):

So D?

OpenStudy (anonymous):

lol u so confused

OpenStudy (kewlgeek555):

i know ^

OpenStudy (anonymous):

why dont you just google it

OpenStudy (ranga):

Yes, D.

OpenStudy (anonymous):

ahhhhhhhahaa u were right nice lol peace

OpenStudy (kewlgeek555):

phew.

OpenStudy (ranga):

It is simple to verify too. Just put n = 1, 2, 3, .... in 4n + 1 and see if you get each term of the sequence.